# LLM-Advisor: An LLM Benchmark for Cost-efficient Path Planning across Multiple Terrains

### v1

@@ Abstract
Multi-terrain cost-efficient path planning is a crucial task in robot navigation, requiring the identification of a path from the start to the goal that not only avoids obstacles but also minimizes travel costs. This is especially crucial for real-world applications where robots need to navigate diverse terrains in outdoor environments, where recharging or refueling is difficult. However, there is very limited research on this topic. In this paper, we develop a prompt-based approach, LLM-Advisor, which leverages large language models (LLMs) as effective advisors for path planning. The LLM-Advisor selectively provides suggestions, demonstrating its ability to recognize when no modifications are necessary. When suggestions are made, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm achieve greater cost efficiency. Since LLM-Advisor may occasionally lack common sense in their suggestions, we propose two hallucination-mitigation strategies. Furthermore, we experimentally verified that GPT-4o performs poorly in zero-shot path planning, even when terrain descriptions are clearly provided, demonstrating its low spatial awareness. We also experimentally demonstrate that using an LLM as an advisor is more effective than directly integrating it into the path-planning loop. Since LLMs may generate hallucinations, using LLMs in the loop of a search-based method (such as A*) may lead to a higher number of failed paths, demonstrating that our proposed LLM-Advisor is a better choice.
@@ I Introduction
Path planning is a task of finding a path from a starting point to a goal while meeting specific conditions such as avoiding obstacles, minimizing travel distance or time, and adhering to other constraints [1]. This task is crucial in multiple fields such as robotics, autonomous navigation, industrial automation, and virtual environment navigation, as it directly impacts the efficiency, safety, and feasibility of operational systems [2]. Many path planning methods have been proposed in recent years, and they can effectively accomplish the planning task while ensuring the validity of the path in most cases [3, 4, 5, 6, 7, 8].
Existing methods for path planning are mainly categorized into search-based and sampling-based approaches. Typical search-based methods include Breadth-First Search (BFS) [9], Depth-First Search (DFS) [10], Best-First Search [11], Dijkstra’s algorithm [12], A* [13], and various A* variants [14, 15, 16, 17, 18, 19, 20]. On the other hand, typical sampling-based methods include Rapidly-exploring Random Tree (RRT) [21], RRT* [22], Informed RRT* [23], RRT-Connect [24], Dynamic RRT [25], Improved RRT [26], etc. However, these methods lack knowledge of the global environment, and they often generate paths with unnecessary detours, which may not be optimal.
@@
Meanwhile, Large Language Models (LLMs) have a powerful capability for understanding global contexts [27, 28, 29, 30], which is especially beneficial in complex problem-solving domains such as path planning [31, 32, 33, 34]. These models have demonstrated the ability of LLMs to handle and reason over long-context inputs, providing global insights that reflect their understanding of the environment, such as identifying obstacles, agents, and the relative positions of goals. However, Ref. [35] found that LLMs perform poorly when dealing with complex, long-term planning and spatial reasoning tasks such as grid-based path planning. LLMs often generate invalid or impractical paths, leading to incomplete routes or collisions, indicating a capability gap when handling detailed spatial complexity. This paper aims to solve the above issues. We claim that by carefully designing the prompts, LLMs can serve as effective advisors for complex path-planning tasks.
In this paper, we first introduce a new task setting: cost-efficient path planning across multiple terrains, which is crucial for unmanned robots. While existing methods mainly consider the travel distance and time, the travel cost is also very important. Robots often operate in environments where recharging or refueling is challenging, such as remote areas, space, or disaster zones. Minimizing energy consumption allows robots to operate for longer periods without the need to recharge, thereby increasing their operational efficacy. Secondly, we propose an innovative path-planning method, LLM-Advisor. LLM-Advisor serves as a post-processing module for the path planning algorithms by leveraging global insights provided by large language models. It suggests improvements to the planned path when the optimal cost-efficient route has not been found. We experimentally found that: 1) By carefully designing the prompts, the proposed LLM-Advisor selectively provides suggestions only when necessary, demonstrating its ability to discern when adjustments are needed and when they are not, and 2) By integrating the proposed hallucination-mitigation strategies, the LLM-Advisor significantly improves its performance in suggesting paths. 3) When suggestions are made, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm prove to be more cost-efficient. The contributions of this paper are threefold:
@@
  • We propose a prompt-based method, LLM-Advisor, that enables LLMs to act as effective advisors after path planning is completed by other methods. Experimental results demonstrate the effectiveness of the proposed prompt and the validity of using LLMs as advisors for path planning.
@@
  • Experimental results demonstrate that LLM-Advisor is capable of selectively making suggestions. Specifically, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm are more cost-efficient.
TABLE I: Terrain difficulty levels, traversal costs, and corresponding colors
@@
Figure 1: A visualization of samples from the created MultiTerraPath dataset
TABLE II: The definitions of traversal costs for the RUGD_v2 dataset
@@
Traversal Cost $\infty$ 1.8 $\infty$ $\infty$ $\infty$ $\infty$ $\infty$ 1.5 1 $\infty$ $\infty$ 3.0
@@
Traversal Cost $\infty$ 2.2 $\infty$ 2.5 $\infty$ $\infty$ $\infty$ $\infty$ 1.2 $\infty$ 3.5 $\infty$
Figure 2: A visualization of samples from the created RUGD_v2 dataset
@@
You are provided with a terrain cost grid description: {terrain_description} You are also provided with the real coordinates of start and goal points, and a description of path list computed by A*. Below is the data for each pair: **Start-End Pair **: - Start: {start} - End: {end} - A* path: {path_description} Analyzing the A* paths, evaluate whether the most cost-efficient path has already been found for each pair. IMPORTANT: **If not, you must suggest a path list, ensure the first point is the exact start coordinates and the last point is the exact end coordinates as provided, display every single coordinate pair**. IMPORTANT: **Make your reply strictly in the following format**: Requirements: If a path is optimal, the answer should be Yes. If a path is not optimal, the answer should be No, the suggested path coordinates (P). The list of coordinates (P) must start at the corresponding start coordinate and end at the corresponding end coordinate. The output format must strictly follow the format below, without any additional information. Output Format: Yes or No, No or [P] IMPORTANT: Only output the real list of coordinates for P. Do not include any extra explanations or text.
Figure 3: The designed prompt for LLM-Advisor
@@
Refer to caption Terrain description: High-cost area with cost 5 is approximately located between grid coordinates (42, 110) and (354, 454). High-cost area with cost 4 is approximately located between grid coordinates (53, 268) and (131, 407). High-cost area with cost 6 is approximately located between grid coordinates (98, 113) and (350, 224). High-cost area with cost 3 is approximately located between grid coordinates (148, 38) and (264, 430). High-cost area with cost 8 is approximately located between grid coordinates (200, 346) and (271, 383). High-cost area with cost 7 is approximately located between grid coordinates (349, 102) and (413, 183). Obstacles are approximately located between grid coordinates (268, 270) and (312, 395). Path description (Detailed): Point 1 at (172, 88) has a terrain cost of 1. Point 2 at (173, 89) has a terrain cost of 1. Point 3 at (174, 90) has a terrain cost of 1. Point 4 at (175, 91) has a terrain cost of 1. Point 5 at (176, 92) has a terrain cost of 1. Point 6 at (177, 93) has a terrain cost of 1. Point 7 at (178, 94) has a terrain cost of 1. Point 8 at (179, 95) has a terrain cost of 1. … Point 285 at (437, 337) has a terrain cost of 5. Point 286 at (438, 337) has a terrain cost of 5. Point 287 at (439, 337) has a terrain cost of 5. Point 288 at (440, 337) has a terrain cost of 5. Point 289 at (441, 337) has a terrain cost of 5. Point 290 at (442, 337) has a terrain cost of 5. Point 291 at (443, 337) has a terrain cost of 5. The total cost of the path is 410.
Figure 4: Examples of terrain descriptions and detailed path descriptions for the A*-planned path in map 1 of the MultiTerraPath dataset
@@
Retrieval-augmented strategy: You are provided with a terrain cost grid description based on historical tasks: {retrieved_terrain_description} Additionally, you have the start and goal coordinates along with computed A-star paths retrieved from similar tasks: {retrieved_path_description} Descriptive output: Requirements: If the path is optimal, respond with “Yes.” If the path is not optimal, respond with “No,” followed by a descriptive sentence indicating that the suggested path is not feasible, and then provide an alternative path. The description should clearly outline the movement from the start coordinate to the end coordinate, step by step. The output must strictly follow the format below, without any additional information. Output Format: Yes or No, No or from $(start_{x},start_{y})$ it goes to $(x_{1},y_{1})$, then $(x_{2},y_{2})$, …, finally, it arrives at $(end_{x},end_{y})$.
Figure 5: Hallucination-mitigation strategies
Figure 6: The A* algorithm search eight neighbors
Figure 7: Visualization of the A* searched paths
@@ II Method
In this section, we first introduce the dataset used in this paper. Then, we describe the details of the proposed LLM-Advisor.
@@ II-A Datasets
MultiTerraPath Dataset. This paper introduces a synthetic dataset for evaluating model performance, MultiTerraPath. Our MultiTerraPath dataset contains 1,000 maps, each sized $500\times 500$. Each map includes several high-cost terrains and one or two impassable obstacles, with one start-to-end pair defined per image. The dataset can be used for tasks such as cost-efficient path planning and obstacle-avoidance path planning. The defined terrains, traversal costs, and their corresponding colors in the images are summarized in Table I. Fig. 1 presents some sample maps from the dataset. The coordinates of the start and end points of each pair are saved in the corresponding JSON file.
RUGD_v2 Dataset. The RUGD dataset [36] supports semantic understanding of unstructured outdoor scenes, including rough and challenging terrains. There are four scenes in the RUGD dataset: creek, park, trail, and village. The size of each image is $688\times 555$. Based on the RUGD dataset, we create one start-to-end pair per image, forming our RUGD_v2 dataset. The traversal costs for different terrains are synthetically defined by GPT-4o (Table II). Fig. 2 shows some examples from the dataset. Note that in real-world applications, it is crucial to conduct thorough experiments to accurately determine the traversal costs for each terrain type. These experiments help to fine-tune the cost values by considering factors such as robot types and the physical properties of the terrain.
@@ II-B LLM-Advisor
The motivations for designing LLM-Advisor are twofold: 1) Existing path planning methods, whether search-based or sampling-based, focus on local regions when conducting path planning. This ensures a reasonable path but may result in unnecessary detours, which significantly affect the total traversal cost. 2) Although some LLM-powered path planning methods have already been developed [37, 34], Ref. [37] leveraged an LLM in an interactive manner, which requires several rounds conversations. Therefore, it is not suitable for real-time operations. Ref. [34] leveraged an LLM to enhance the efficiency of the A* algorithm by generating a candidate list during the search process. However, the LLM-A* approach [34] led to an increased rate of failures in finding valid paths. Considering the above two issues, we propose using LLMs as advisors as a viable solution. However, due to the LLM’s low spatial awareness and tendency for hallucinations, careful prompt design and the implementation of mitigation strategies are essential.
Main Prompt. Prompt design is a key component of the proposed LLM-Advisor. Specifically, when designing the main prompt, the terrain description is included as knowledge, as it provides critical information about the path planning environment. To address the issue of low spatial awareness, we also incorporate a detailed description of the planned path. By asking the LLMs to evaluate both the terrain and the path, we aim to enhance their spatial awareness and improve its performance. The code in Fig. 3 shows the designed prompt, where the GPT-4o is utilized. Examples of the terrain descriptions and detailed path descriptions used are presented in Fig. 4.
Hallucination-mitigation Strategies. Furthermore, we propose two mitigation strategies to alleviate the hallucinations of LLM-Advisor. Details are given in Fig. 5.
  • DescPath. Since LLM-Advisor relies heavily on the contextual understanding capabilities of LLMs, directly generating coordinate points of the path may introduce hallucinations. Therefore, we propose generating a descriptive path (DescPath) instead of just listing individual path coordinates.

### v2

@@ Abstract
Cost-efficient path planning across multiple terrains is a crucial task in robot navigation, requiring the identification of a path from the start to the goal that not only avoids obstacles but also minimizes the overall travel cost. This is especially crucial for real-world applications where robots need to navigate diverse terrains in outdoor environments with limited opportunities for recharging or refueling. Despite its practical importance, cost-efficient path planning across heterogeneous terrains has received relatively limited attention in prior work. In this paper, we propose LLM-Advisor, a prompt-based, planner-agnostic framework that leverages large language models (LLMs) as non-decisive post-processing advisors for cost refinement, without modifying the underlying planner. While we observe that LLMs may occasionally produce implausible suggestions, we introduce two effective hallucination-mitigation strategies. We further introduce two datasets, MultiTerraPath and RUGD_v2, for systematic evaluation of cost-efficient path planning. Extensive experiments reveal that state-of-the-art LLMs, including GPT-4o, GPT-4-turbo, Gemini-2.5-Flash, and Claude-Opus-4, perform poorly in zero-shot terrain-aware path planning, highlighting their limited spatial reasoning capability. In contrast, the proposed LLM-Advisor (with GPT-4o) improves cost efficiency for 72.37% of A*-planned paths, 69.47% of RRT*-planned paths, and 78.70% of LLM-A*-planned paths. On the MultiTerraPath dataset, LLM-Advisor demonstrates stronger performance on the hard subset, further validating its applicability to real-world scenarios.
@@ I Introduction
Path planning is a task of finding a path from a starting point to a goal while meeting specific conditions such as avoiding obstacles, minimizing travel distance or time, and meeting other constraints [31, 23, 24, 11, 28]. This task is crucial in multiple fields such as robotics, autonomous navigation, industrial automation, and virtual environment navigation, as it directly impacts the operational efficiency, safety, and feasibility of systems [6]. Many path planning methods have been proposed in recent years, and they can effectively accomplish the planning task while ensuring the validity of the path in most cases [13, 3, 18, 12, 5].
Existing methods are mainly search-based and sampling-based approaches. Typical search-based methods include Breadth-First Search (BFS) [30], Depth-First Search (DFS) [15], Best-First Search [19], Dijkstra’s algorithm [14], A* [4], and various A* variants [17, 8]. On the other hand, typical sampling-based methods include Rapidly-exploring Random Tree (RRT) [10], RRT* [7], Informed RRT* [2], RRT-Connect [9], Dynamic RRT [29], Improved RRT [22], etc. Despite their effectiveness, these planners perform global optimization over discretized graph representations, whose structure and connectivity are constrained by map resolution, neighborhood definitions, and hand-crafted cost models. For search-based planners, heuristic-guided expansions operate on predefined graph neighborhoods, which limits the expressiveness of the search space and its ability to capture large-scale structural patterns or long-range cost trends present in the continuous environment. As a result, even with access to the full cost map, the resulting paths may still exhibit unnecessary detours or suboptimal cost trade-offs.
@@
Meanwhile, Large Language Models (LLMs) have strong capabilities for understanding global contexts [27], which is particularly beneficial in complex problem-solving domains such as path planning [21, 20, 16]. These models have demonstrated the ability of LLMs to handle and reason over long-context inputs, providing global insights that reflect their understanding of the environment, such as identifying obstacles, agents, and the relative positions of goals. However, Aghzal et al. [1] found that LLMs perform poorly when dealing with complex, long-term planning and spatial reasoning tasks such as grid-based path planning. LLMs often generate invalid or impractical paths, leading to incomplete routes or collisions, indicating a capability gap when handling detailed spatial complexity. This paper aims to address these issues. We claim that by carefully designing the prompts, LLMs can serve as effective advisors for complex path-planning tasks.
In this paper, we introduce a new task: cost-efficient path planning across multiple terrains, a critical challenge for unmanned robots. Unlike existing methods that mainly optimize distance or time, we emphasize travel cost. Given a path represented as a sequence of connected segments, the total travel cost is defined as the sum over all path segments, where each segment cost is computed as the product of its Euclidean length and the corresponding terrain cost. This formulation is particularly relevant for robots operating in environments where recharging is difficult (e.g., remote areas, space, and disaster zones). Reducing energy use directly extends operational time and effectiveness.
However, achieving cost-efficient navigation remains challenging in practice. Although classical planners such as A* are globally optimal with respect to a given discrete graph formulation, in practical robot navigation settings they typically operate on discretized maps with limited connectivity, due to sensing resolution, computational constraints, or map availability. As a result of discretization effects and limited connectivity, the resulting paths may still be suboptimal with respect to the underlying continuous terrain and long-range cost structure. This motivates the need for an additional mechanism that can reason over the global terrain layout and propose path refinements that are not representable within the planner’s original discrete search space, while leaving A*’s optimality guarantees on the discrete graph intact. Specifically, in our experimental setting, A* operates on a discretized grid graph with fixed 8-connected connectivity and uses the Manhattan distance to the goal as the heuristic function.
We then propose LLM-Advisor, a post-processing module that leverages large language models to refine paths generated by classical planners. It selectively suggests improvements only when needed and incorporates hallucination-mitigation strategies, achieving significant gains: 72.37% improved for A*, 69.47% for RRT*, and 78.70% for LLM-A*. Unlike prior LLM approaches that rely on end-to-end generation or interactive queries, our method uses LLMs in an advisory role. With careful prompt design and hallucination control, LLM-Advisor improves cost efficiency without retraining or modifying the planner. Importantly, we do not aim to outperform optimally designed continuous or high-connectivity planners; instead, we target practical scenarios with fixed discretization, limited connectivity, or legacy implementations, where LLM-Advisor provides a planner-agnostic and low-risk mechanism for cost refinement. In summary, the contributions of this paper are threefold:
@@
  • We propose a prompt-based method, LLM-Advisor, that enables LLMs to act as effective advisors after path planning is completed by other methods.
@@
  • Experimental results demonstrate that LLM-Advisor is capable of selectively making suggestions. Specifically, 72.37% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm are more cost-efficient. The proposed hallucination-mitigation strategies significantly reduce hallucinations.
TABLE I: Terrain difficulty levels, travel costs, and corresponding colors.
@@
Figure 1: MultiTerraPath samples. Easy maps contain separated convex high-cost regions, while hard maps include irregular, connected terrains with no purely zero-cost paths, requiring cost-aware trade-offs.
Figure 2: A visualization of samples from the created RUGD_v2 dataset.
TABLE II: The definitions of travel costs for the RUGD_v2 dataset.
@@
Travel Cost $\infty$ 1.8 $\infty$ $\infty$ $\infty$ $\infty$ $\infty$ 1.5 1 $\infty$ $\infty$ 3.0
@@
Travel Cost $\infty$ 2.2 $\infty$ 2.5 $\infty$ $\infty$ $\infty$ $\infty$ 1.2 $\infty$ 3.5 $\infty$
Figure 3: The pipeline of proposed LLM-Advisor.
@@
You are provided with a terrain cost grid description: {terrain_description}. You are also provided with the real coordinates of the start and goal points, and a description of the path list computed by {existing path planning method}. Below is the data for each pair: **Start-End Pair **: - Start: {start} - End: {end} - {existing path planning method}-planned path: {path_description} Analyzing the {existing path planning method}-planned path, evaluate whether the most cost-efficient path has already been found for each pair. IMPORTANT: **If not, you must suggest a path list, ensure the first point is the exact start coordinates and the last point is the exact end coordinates as provided, display every single coordinate pair**. IMPORTANT: **Make your reply strictly in the following format**: Requirements: If a path is optimal, the answer should be Yes. If a path is not optimal, the answer should be No, the suggested path coordinates (P). The list of coordinates (P) must start at the corresponding start coordinate and end at the corresponding end coordinate. The output format must strictly follow the format below, without any additional information. Output Format: Yes or No, No or [P] IMPORTANT: Only output the real list of coordinates for P. Do not include any extra explanations or text.
Figure 4: The designed prompt for LLM-Advisor.
@@
Refer to caption Terrain description: High-cost area with cost 5 is approximately located between grid coordinates (42, 110) and (354, 454). High-cost area with cost 4 is approximately located between grid coordinates (53, 268) and (131, 407)… Obstacles are approximately located between grid coordinates (268, 270) and (312, 395). Path description (Detailed): Point 1 at (172, 88) has a terrain cost of 1. Point 2 at (173, 89) has a terrain cost of 1. Point 3 at (174, 90) has a terrain cost of 1 … Point 290 at (442, 337) has a terrain cost of 5. Point 291 at (443, 337) has a terrain cost of 5. The total cost of the path is 410.
Figure 5: Examples of terrain descriptions and detailed path descriptions for the A*-planned path in map 1 of the MultiTerraPath dataset. The terrain descriptions are automatically generated by grouping grid cells with similar cost values, extracting their bounding boxes, and converting them into concise text summaries. This process is fully automatic and scales linearly with map size.
@@
Retrieval-augmented strategy: You are provided with a terrain cost grid description based on historical tasks: {retrieved_terrain_description}. Additionally, you have the start and goal coordinates along with the computed {existing path planning method}-planned path retrieved from similar tasks: {retrieved_path_description}. Descriptive output: Requirements: If the path is optimal, respond with “Yes.” If the path is not optimal, respond with “No,” followed by a descriptive sentence indicating that the suggested path is not feasible, and then provide an alternative path. The description should clearly outline the movement from the start coordinate to the end coordinate, step by step. The output must strictly follow the format below, without any additional information. Output Format: Yes or No, No or from $(start_{x},start_{y})$ it goes to $(x_{1},y_{1})$, then $(x_{2},y_{2})$, …, finally, it arrives at $(end_{x},end_{y})$.
Figure 6: Hallucination-mitigation strategies.
@@ II-A Datasets
MultiTerraPath Dataset. This paper introduces MultiTerraPath, a synthetic dataset designed for evaluating model performance in cost-aware path planning. MultiTerraPath contains 2,000 maps, each with a resolution of $500\times 500$, comprising an easy subset with 1,000 maps and a hard subset with 1,000 maps. Easy map contains convex high-cost regions that are well separated by zero-cost areas, allowing straightforward low-cost routing. Hard map, in contrast, features irregular and concave terrain regions (e.g., corridor- or ribbon-shaped structures), allow terrain regions to be connected or overlapping to better mimic real-world continuous landscapes, and include cases where no purely zero-cost path exists, thereby forcing explicit cost-aware trade-offs. The dataset can be used for tasks such as cost-efficient path planning and obstacle-avoidance path planning. The defined terrains, travel costs, and their corresponding colors in the images are summarized in Table I. Fig. 1 presents some sample maps from the dataset. The coordinates of the start and end points of each pair are saved in the corresponding JSON file.
RUGD_v2 Dataset. The RUGD dataset [25] provides pixel-level semantic annotations for unstructured outdoor environments, covering four scenes: creek, park, trail, and village, with image resolution $688\times 555$. We create one start–goal pair per image to construct our RUGD_v2 dataset. Terrain-dependent travel costs are synthetically assigned using GPT-4o (Table II), and regions with infinite cost are treated as impassable obstacles. Fig. 2 shows sample images. In real applications, terrain costs should be calibrated through physical experiments considering robot type and terrain properties.
@@ II-B LLM-Advisor
The motivations for designing LLM-Advisor are twofold: 1) Existing path planning methods, whether search-based or sampling-based, operate on discretized maps with limited connectivity, due to sensing resolution, computational constraints, or map availability. While this often yields feasible paths, it can lead to unnecessary detours, increasing the total travel cost. 2) Although some LLM-powered path planning methods have already been developed [26, 16], Ref. [26] leveraged an LLM in an interactive manner, which requires multiple rounds of conversation. Therefore, it is not suitable for real-time operations. Ref. [16] leveraged an LLM to enhance the efficiency of the A* algorithm by generating a candidate list during the search process. However, the LLM-A* approach [16] exhibited a higher failure rate in finding valid paths. To address these two issues, we propose using LLMs as non-decisive advisors. However, due to the limited spatial awareness and tendency for hallucinations in LLMs, careful prompt design and the implementation of mitigation strategies are essential. The overall pipeline of LLM-Advisor is provided in Fig. 3.
Main Prompt. Prompt design is a key component of the proposed LLM-Advisor. Specifically, the main prompt incorporates terrain descriptions as prior knowledge, providing critical information about the path planning environment. To address the issue of limited spatial awareness, we also incorporate a detailed description of the planned path. By prompting the LLMs to evaluate both the terrain and the path, we aim to enhance their spatial awareness and overall performance. The code in Fig. 4 shows the designed prompt, where the GPT-4o is utilized. Examples of the terrain descriptions and detailed path descriptions used are presented in Fig. 5. To ensure generalizability, our prompt design follows a modular template that separates terrain encoding, path description, and evaluation directive. This structure enables easy adaptation across tasks and environments. Moreover, the same prompt format was used successfully on both the synthetic MultiTerraPath and real-scene RUGD v2 datasets, suggesting that the prompt structure generalizes well without manual tuning per environment.
Hallucination-mitigation Strategies. Furthermore, we propose two strategies to alleviate the hallucinations in LLM-Advisor. Details are given in Fig. 6.
  • DescPath. Since LLM-Advisor relies heavily on the contextual reasoning capabilities of LLMs, directly generating coordinate points of the path may introduce hallucinations. Therefore, we propose generating a descriptive path (DescPath) instead of just listing individual path coordinates.